OpenStudy (anonymous):

\[T=2\Pi \sqrt{L/32}\] for L

OpenStudy (blockcolder):

\[\frac{T}{2\pi}=\sqrt{\frac{L}{32}}\\ \frac{T^2}{4\pi^2}=\frac{L}{32}\\ L=\frac{8T^2}{\pi^2}\]